OpenStudy (anonymous):

solve this (5-z)(3z+8)=0

OpenStudy (imstuck):

What you do here is set each parenthesis equal to 0 and solve for z. The first one, then, would be 5 - z = 0...-z = -5...or z = 5. The next one would be 3z + 8 = 0...3z = -8...z = -8/3

OpenStudy (anonymous):

i got -8/3

OpenStudy (imstuck):

Very good! Those are your solutions to the original quadratic equation (which they didn't give you in its standard form; they already factored it for you).

OpenStudy (anonymous):

\[-3 z^2+7 z+40 = 0\] this is a quadratic equation which gives to solutions -8/3 and 5.